# Enhanced Electromagnetic Memory

### v1

@@ Abstract
We show that the amplitude of electromagnetic memory can be significantly enhanced in comparison to known estimates. In a Lorentz breaking phase of lowered phase velocity of light, there exist critical spacetime directions of memory-source emission along the effective light cone, about which the total memory offset receives order of magnitude increases. The same amplification is already present in the Lorentz preserving case by considering ultra-relativistic memory-source charges. These observations may pave the way for a first observation of the phenomenon of memory and laboratory tests of the concepts of asymptotic symmetries and soft theorems.
@@ I Introduction
Gravitational radiation not only consists of the observed gravitational waves, but always also includes a non-wave-like component that permanently alters the background structure of asymptotically flat spacetime in the form of a net shift in proper distances. This phenomenon is known as gravitational memory [1, 2, 3, 4, 5, 6, 7, 8] (see also [9, 10, 11, 12, 13, 14, 15]) and represents one of the fundamental predictions of general relativity that has not yet been observed. This additional contribution to the radiation can be viewed as a back-reaction effect of emitted radiative energy-momentum. The resulting memory-induced transition on the state of the background spacetime is radically related to a supertranslation of asymptotic Bondi-Metzner-Sachs (BMS) symmetries [16, 17, 6, 18, 11, 19] and can therefore be viewed as a spontaneous breaking of the latter. Moreover, as a “net” background effect, memory probes the zero frequency limit of the radiating gauge field, such that the phenomenon is equally connected to the gravitational soft theorems [20, 11].
The relation to soft factors, as well as the fact that gravitational memory can be viewed as a mere consequence of the wave equation in a radiative process [8, 14, 21, 22, 15], suggests that an analogous phenomenon must exist within the theory of electromagnetism. Indeed, based on the gravitational knowledge, it was possible to describe an electromagnetic memory [23] that manifests itself as a non-zero offset within the radiative transverse vector potential, representing the classical manifestation of photon soft theorems and the asymptotic $U(1)$ symmetry of electric charge conservation [24, 25]. In contrast to the gravitational case, however, electromagnetic waves themselves do not carry any corresponding charge, such that a production of electromagnetic memory necessarily requires an emission of unbound charged matter.
@@ II.1 Dynamical Memory Equation
$-\frac{1}{v_{\text{em}}^{2}}\ddot{A}^{\alpha}+\,\partial^{2}A^{\alpha}=-\mu J^{\alpha}\,.$ (1)
@@
Let’s assume the idealized model of a fast electric charge $q$ that is instantly emitted from a localized process at the coordinate origin with constant 3-velocity $\frac{d\mathbf{x}_{q}}{dt}=v_{q}\hat{\mathbf{n}}_{q}$ in a given direction characterized by a unit radial outward vector $\hat{\mathbf{n}}_{q}=\hat{\mathbf{n}}(\Omega_{q})$ defined through the angles $\Omega_{q}=\{\theta_{q},\phi_{q}\}$ in spherical coordinates [Eq. (27)]. The associated 4-velocity reads $U_{q}^{\alpha}=\frac{dx_{q}^{\alpha}(\tau)}{d\tau}=\gamma_{q}V^{\alpha}_{q}$, where $V^{\alpha}_{q}\equiv(c,v_{q}\hat{\mathbf{n}}_{q})$ and $\gamma_{q}\equiv 1/\sqrt{1-(v_{q}/c)^{2}}$ and the 4-current density becomes
[…]
@@ II.2 A General Electromagnetic Memory Solution
Depending on the sign of the Green’s function factor $\mathcal{V}$, one obtains two different solutions.
@@ II.3 Directions of Memory Enhancement
$\Delta A^{T}(r,\theta,\theta_{q}=0)$ =$=\begin{cases}\frac{q\mu v_{q}}{4\pi r}\frac{\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta<1$},\\ \frac{q\mu v_{q}}{4\pi r}\frac{-\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta>1$},\end{cases}$ (20)
@@
Figure 2: The total electromagnetic memory amplitude $\Delta A^{T}(1,\theta,\theta_{q}=0)$ [Eq. (20)] in SI units of [Vs/m] for a unit charge $q=1$[C] and distance $r=1$[m], as a function of the azimuthal angle $\theta$ and for different values of charge emission speeds $v_{q}$. For concreteness, the phase velocity of light in the medium is set to $v_{\text{em}}=0.9c$, with $\mu=\mu_{0}$ and $\epsilon\approx 1.23\,\epsilon_{0}$. For a superluminal emission of the charge $v_{q}>v_{\text{em}}$ (solid blue) there exists an unprotected direction of unbounded memory sourcing $\theta_{c}=\arccos[v_{\text{em}}/v_{q}]=0.2089$, indicated by a black dashed vertical line, around which the amplitude of electromagnetic memory is considerably enhanced. If $v_{q}=v_{\text{em}}$ (dotdashed orange) the critical angle coincides with the direction of emission of the source $\hat{\mathbf{n}}=\hat{\mathbf{n}}_{q}$. For a subluminal emission velocity $v_{q}<v_{\text{em}}$ (dashed green and dotted red), the electromagnetic memory offset remains finite for any values of $\theta$, although the amplitude is still significantly increased at low values of the angle $\theta$ as soon as $v_{q}\gtrsim 0.99\,v_{\text{em}}$.
@@
While the divergence in the memory amplitude at the exact critical direction in Eq. (17) is unphysical - it directly translates into a divergence in the emitted radiative energy as shown in appendix B - and the underlying assumptions of memory computations break down, the significant increase in memory offset in the vicinity of the critical directions is real [30]. This is also shown by the fact that the enhancement already occurs in an entirely finite and well-defined manner within a Lorentz-preserving vacuum phase, as discussed above. Furthermore, we explicitly show in appendix C that the divergence condition can geometrically be explained as arising whenever the source charge is asymptotically emitted along the spacetime direction of a past null cone characterized by the phase velocity of light $v_{\text{em}}$. Such critical directions will therefore only be present at luminal or superluminal speeds $v_{q}$ of charge emission. We note that Eq. (17) is equivalent to the condition on the emission angle of Cherenkov radiation [32, 33, 34]  [^2] The crucial difference to Cherenkov radiation is that the actual production of memory radiation only occurs during the acceleration process of the source.
@@ III Observational Implications
that reveals the observational implications of a non-zero offset within the radiative vector potential $\Delta A_{T}^{i}\neq 0$. Assuming a finite burst of electromagnetic radiation, an integration of this equation results in
[…]
@@
Hence, once all electromagnetic radiation has passed, the test charge $Q$ receives a non-zero transverse velocity kick [23]. This memory kick is distinguished from radiation pressure associated to the Poynting vector through its transversality and its linearity with respec to $Q$.
@@
This is however not the only signature of a direct measurement of electromagnetic memory. In a realistic experiment, the emission of a charge (or scattering) takes place over a characteristic time scale that we will denote as $\Delta u$, at which the velocities of outgoing (and incoming) charges depart from their asymptotic constant value. Such a time scale is equivalently associated to a “size” of the emitted charges and effectively smoothens out the delta function in asymptotic retarded time within Eq. (14). This in turn results in a characteristic rise-time-scale of memory production by altering the Heaviside stepfunciton within the solutions in Eqs. (15) and (16). The corresponding transverse electric field in the radiation zone associated to the radiation of electromagnetic memory in Eq. (23) will consequently be of a typical dome shape with obvious non-zero time integral, representing another characteristic signature of electromagnetic memory to look for in realistic observations [29].
@@
Different experiment types have already been proposed for a direct measurement of electromagnetic memory, both in macroscopic and microscopic setups of memory production. One concrete proposal [29] suggests the use of a dipole antenna generating two opposite unbound charge pulses that induce an emission of electromagnetic memory  [^3] Note that for a simultaneous emission of multiple current pulses in different directions, the results in section II can simply be stacked. Therefore, an emission of opposite currents in a dipole antenna, for instance, would result in a memory production, in which the features of Fig. 2 are mirrored at $\theta=\pi/2$ with opposite signs., which can be observed as a net current within a second receiver antenna, as an indication for a non-zero time integral of the radiative electric field. Another thought experiment [28] considers an “explosion” of charged-particles from a localized source surrounded by a sphere of superconducting nodes, within which the memory effect could be observed through a permanent imprint within the relative phases of the nodes.
With this work, we show that in similar experimental setups, the amplitude of electromagnetic memory can significantly be enhanced  [^4] To get a feeling for the values $\Delta A^{T}\sim 10^{2}-10^{4}$[Vs/m] in SI units presented in Fig. 2, consider the following estimates: At length scales of meters, the time scales of the experiment will be of nanoseconds, such that a typical current of $10^{-3}$[A] carries a charge of order $q\sim 10^{-12}$[C]. With these numbers, a two-orders-of-magnitude gain in the potential amplitude lifts the expected strength of the electric field from $E_{\text{max}}\sim\frac{q}{[\text{C}]}\frac{100}{[\text{ns}]}\,[\text{Vs/m}]\sim 10^{-1}$ [V/m] to $10$ [V/m]. within a Lorentz breaking phase, whenever the superluminal regime $v_{q}\geq v_{\text{em}}$ is reached. If the speed of charge emission $v_{q}$ exceeds the effective speed of light $v_{\text{em}}$, the single critical direction unfolds to a circle at a characteristic angle of
$\theta=\arccos(v_{\text{em}}/v_{q})\,,$ (24)
@@
about the direction of charge propagation. As already mentioned, unlike in the gravitational analogue, the structure of electromagnetic memory is such that an important increase in amplitude can already be achieved in a Lorentz preserving situation as proposed in [29, 28], when the speed of charge emission approaches the vacuum speed of light $c$. However, both the speed of a current pulse in a dipole antenna, as well as the emission velocity of charged particles for instance in $\beta$-decay  [^5] The idealized situation of a charge emitted at constant velocity $v_{q}$ considered in this work is indeed an excellent model of $\beta$-decay [40, 41]. are not expected to reach the threshold of enhancement of $v_{q}>0.99\,c$ identified in section II.3. In this context, a down-scale of the effective phase velocity of light through Lorentz violating effects is crucial to reach the regime of significant memory amplification.
@@
In general, a decrease of the phase velocity of light can either be realized within particular media, as explicitly presented in section II in the case of a simple dielectric medium, or in the vacuum within adequate slow-wave structures, such as traveling-wave tubes (TWT). Of course, the conception of a realistic experimental setup to measure electromagnetic memory through the Lorentz breaking enhancement identified in this work still requires at least two significant steps that we want to address in future work. The first is a refined, quantitative computation of the memory amplitude within the given situation, taking into account finite-size- as well as possible shock-wave-effects that cannot be captured through the Green’s function of the wave equation  [^6] The computational techniques developed in the context of Cherenkov radiation might be of great use.. The second is an elaboration on the measurement technique, where it must in particular be ensured that the measurement time-scale is quick enough to capture the memory signal within the unbounded state of the emitted charges and that the memory kick can be discriminated from possible Poynting vector kicks associated with radiation pressure.
@@ IV Conclusion
While the latter case necessarily requires a reduction of the effective phase velocity of light $v_{\text{em}}<c$, the former can also be achieved at a vacuum speed of light $v_{\text{em}}=c$ with the emission of ultra-relativistic charges of $v_{q}\gtrsim 0.99c$.
@@
Although the actual divergence in the memory amplitude as well as the emitted radiative energy remains intangible with the treatment of this work, the enhancement in the vicinity of the critical directions is physical, as shown by the presence of the entirely well-defined effect in the subluminal vacuum case. An important future task will be to refine the analysis to compute the realistic peak in memory offset increase in specific local Lorentz breaking setups and conceptualize concrete measurement strategies of the electromagnetic memory amplitude. Moreover, it might be interesting to explore a connection of the results with known superluminal and supersonic phenomena.

### v2

# Enhancement of Electromagnetic Memory Effects

@@ Abstract
We show that the amplitude of electromagnetic memory can be significantly enhanced in comparison to known estimates. In a Lorentz breaking phase of lowered phase velocity of light, there exist critical spacetime directions of memory-source emission along the effective light cone, about which the total memory offset receives order of magnitude increases. The same amplification is already present in the Lorentz preserving case by considering ultrarelativistic memory-source charges. These observations may pave the way for a first observation of the phenomenon of memory and laboratory tests of the concepts of asymptotic symmetries and soft theorems.
@@ I Introduction
Gravitational radiation not only consists of the observed gravitational waves, but always also includes a nonwavelike component that permanently alters the background structure of asymptotically flat spacetime in the form of a net shift in proper distances. This phenomenon is known as gravitational memory [1, 2, 3, 4, 5, 6, 7, 8] (see also [9, 10, 11, 12, 13, 14, 15]) and represents one of the fundamental predictions of general relativity that has not yet been observed. This additional contribution to the radiation can be viewed as a back-reaction effect of emitted radiative energy momentum. The resulting memory-induced transition on the state of the background spacetime is radically related to a supertranslation of asymptotic Bondi-Metzner-Sachs (BMS) symmetries [16, 17, 6, 18, 11, 19] and can therefore be viewed as a spontaneous breaking of the latter. Moreover, as a “net” background effect, memory probes the zero frequency limit of the radiating gauge field, such that the phenomenon is equally connected to the gravitational soft theorems [20, 11, 21].
The relation to soft factors, as well as the fact that gravitational memory can be viewed as a mere consequence of the wave equation in a radiative process [8, 14, 22, 23, 15], suggests that an analogous phenomenon must exist within the theory of electromagnetism. Indeed, based on the gravitational knowledge, it was possible to describe an electromagnetic memory [24] that manifests itself as a nonzero offset within the radiative transverse vector potential, representing the classical manifestation of photon soft theorems and the asymptotic $U(1)$ symmetry of electric charge conservation [25, 26, 27, 28, 21]. In contrast to the gravitational case, however, electromagnetic waves themselves do not carry any corresponding charge, such that a production of electromagnetic memory necessarily requires an emission of unbound charged matter.
@@ II.1 Dynamical Memory Equation
$\frac{1}{v_{\text{em}}^{2}}\ddot{A}^{\alpha}+\,\partial^{2}A^{\alpha}=-\mu J^{\alpha}\,.$ (1)
@@
Let’s assume the idealized model of a fast electric charge $q$ that is instantly emitted from a localized process at the coordinate origin with constant 3-velocity $\frac{d\mathbf{x}_{q}}{dt}=v_{q}\hat{\mathbf{n}}_{q}$ in a given direction characterized by a unit radial outward vector $\hat{\mathbf{n}}_{q}=\hat{\mathbf{n}}(\Omega_{q})$ defined through the angles $\Omega_{q}=\{\theta_{q},\phi_{q}\}$ in spherical coordinates [Eq. (25c)]. The associated 4-velocity reads $U_{q}^{\alpha}=\frac{dx_{q}^{\alpha}(\tau)}{d\tau}=\gamma_{q}V^{\alpha}_{q}$, where $V^{\alpha}_{q}\equiv(v_{\text{em}},v_{q}\hat{\mathbf{n}}_{q})$ and $\gamma_{q}\equiv 1/\sqrt{1-(v_{q}/c)^{2}}$ and the 4-current density becomes
$J_{q}^{\alpha}(t,\vec{x})=q\int d\tau\,U_{q}^{\alpha}\,\delta^{4}(x-x_{q}(\tau))=q\,V_{q}^{\alpha}\delta^{3}(\mathbf{x}-\mathbf{x}_{q}(t))\,.$ (4)
@@ II.2 A General Electromagnetic Memory Solution
Depending on the sign of the Green’s function factor $\mathcal{V}$ in Eq. (12), one obtains two different solutions.
@@ II.3 Directions of Memory Enhancement
$\Delta A^{T}(r,\theta,\theta_{q}=0)$ $=\begin{cases}\frac{q\mu v_{q}}{4\pi r}\frac{\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta<1$},\\ \frac{q\mu v_{q}}{4\pi r}\frac{-\sin\theta}{1-\frac{v_{q}}{v_{\text{em}}}\cos\theta},&\text{if $\frac{v_{q}}{v_{\text{em}}}\cos\theta>1$},\end{cases}$ (20)
@@
Figure 2: The total electromagnetic memory amplitude $\Delta A^{T}(1,\theta,\theta_{q}=0)$ [Eq. (20)] in SI units of [Vs/m] for a unit charge q=1q=1[C] and distance r=1r=1[m], as a function of the azimuthal angle $\theta$ and for different values of charge emission speeds $v_{q}$. For concreteness, the phase velocity of light in the medium is set to $v_{\text{em}}=0.9c$, with $\mu=\mu_{0}$ and $\epsilon\approx 1.23\,\epsilon_{0}$. For a superluminal emission of the charge $v_{q}>v_{\text{em}}$ (solid blue) there exists an unprotected direction of unbounded memory sourcing $\theta_{c}=\arccos[v_{\text{em}}/v_{q}]=0.2089$, indicated by a black dashed vertical line, around which the amplitude of electromagnetic memory is considerably enhanced. If $v_{q}=v_{\text{em}}$ (dotdashed orange) the critical angle coincides with the direction of emission of the source $\hat{\mathbf{n}}=\hat{\mathbf{n}}_{q}$. For a subluminal emission velocity $v_{q}<v_{\text{em}}$ (dashed green and dotted red), the electromagnetic memory offset remains finite for any values of $\theta$, although the amplitude is still significantly increased at low values of the angle $\theta$ as soon as $v_{q}\gtrsim 0.99\,v_{\text{em}}$.
@@
While the divergence in the memory amplitude at the exact critical direction in Eq. (17) is unphysical - it directly translates into a divergence in the emitted radiative energy as shown in appendix B and the underlying assumptions of memory computations break down - the significant increase in memory offset in the vicinity of the critical directions is real [33]. This is also shown by the fact that the enhancement already occurs in an entirely finite and well-defined manner within a Lorentz-preserving vacuum phase, as discussed above. Furthermore, we explicitly show in appendix C that the divergence condition can geometrically be explained as arising whenever the source charge is asymptotically emitted along the spacetime direction of a past null cone characterized by the phase velocity of light $v_{\text{em}}$. Such critical directions will therefore only be present at luminal or superluminal speeds $v_{q}$ of charge emission. We note that Eq. (17) is equivalent to the condition on the emission angle of Cherenkov radiation [35, 36, 37]  [^2] The crucial difference to Cherenkov radiation is that the actual production of memory radiation only occurs during the acceleration process of the source. The above arguments strongly suggest that the enhanced memory burst is a shock-wave phenomenon, generally characterized by the presence of a formally singular potential.
@@ III Observational Implications
that reveals the observational implications of a nonzero offset within the radiative vector potential $\Delta A_{T}^{i}\neq 0$. Assuming a finite burst of electromagnetic radiation, an integration of this equation results in
$\Delta v^{i}=\frac{Q}{M}\int_{-\infty}^{\infty}dtE^{i}=-\frac{Q}{M}\Delta A_{T}^{i}\,$, (22)
@@
Hence, once all electromagnetic radiation has passed, the test charge $Q$ receives a nonzero transverse velocity kick [24] [^3] To prevent confusion we want to briefly mention here that the concern about the non-existence of the velocity kick on asymptotic test charges expressed in Ref. [54] is based on a misunderstanding: While that reference computes the Coulombic field of outgoing free charged particles, and therefore correctly infers that in such a physical situation there is no energy-carrying radiation, the electromagnetic memory effect is a locally produced radiation field that requires the acceleration or scattering of unbound charges.. This memory kick is distinguished from radiation pressure associated to the Poynting vector through its transversality and its linearity with respec to $Q$.
@@
This is however not the only signature of a direct measurement of electromagnetic memory. In a realistic experiment, the emission of a charge (or scattering) takes place over a characteristic time scale that we will denote as $\Delta u$, at which the velocities of outgoing (and incoming) charges depart from their asymptotic constant value. Such a time scale is equivalently associated to a “size” of the emitted charges and effectively smoothens out the delta function in asymptotic retarded time within Eq. (14). This in turn results in a characteristic rise-time-scale of memory production by altering the Heaviside stepfunciton within the solutions in Eqs. (15) and (16). The corresponding transverse electric field in the radiation zone in Eq. (23) will consequently be of a typical dome shape with obvious non-zero time integral, representing another characteristic signature of electromagnetic memory to look for in realistic observations [32].
@@
Different experiment types have already been proposed for a direct measurement of electromagnetic memory, both in macroscopic and microscopic setups of memory production. One concrete proposal [32] suggests the use of a dipole antenna generating two opposite unbound charge pulses that induce an emission of electromagnetic memory  [^4] Note that for a simultaneous emission of multiple current pulses in different directions, the results in section II can simply be stacked. Therefore, an emission of opposite currents in a dipole antenna, for instance, would result in a memory production, in which the features of Fig. 2 are mirrored at $\theta=\pi/2$ with opposite signs.. This signal could then be observed as a net current within a second receiver antenna, as an indication for a non-zero time integral of the radiative electric field. Another thought experiment [31] considers an “explosion” of charged-particles from a localized source surrounded by a sphere of superconducting nodes, within which the memory effect could be observed through a permanent imprint within the relative phases of the nodes.
With this work, we show that in similar experimental setups, the amplitude of electromagnetic memory can significantly be enhanced within a Lorentz breaking phase, whenever the superluminal regime $v_{q}\geq v_{\text{em}}$ is reached, forming a circle of amplitude enhancement about the direction of charge propagation at the characteristic angle
$\theta=\arccos(v_{\text{em}}/v_{q})\,.$ (24)
@@
As already mentioned, unlike in the gravitational analog, the structure of electromagnetic memory is such that an important increase in amplitude can already be achieved in a Lorentz preserving situation as proposed in [32, 31], when the speed of charge emission approaches the vacuum speed of light $c$. However, both the speed of a current pulse in a dipole antenna, as well as the emission velocity of charged particles for instance in $\beta$-decay  [^5] The idealized situation of a charge emitted at constant velocity $v_{q}$ considered in this work is indeed an excellent model of $\beta$-decay [50, 51]. are not expected to reach the threshold of enhancement of $v_{q}>0.99\,c$ identified in section II.3. In this context, a down scale of the effective phase velocity of light through Lorentz violating effects is crucial to reach the regime of significant memory amplification.
To get a feeling for the significance of the enhancement, consider the values $\Delta A^{T}\sim\frac{q}{\text{[C]}}\frac{\text{[m]}}{r}10^{2}-10^{4}$[Vs/m] in SI units presented in Fig. 2, as a function of charge and distance. Both for the dipole antenna proposal in [32] and typical $1$ MeV electron beams, a somewhat conservative value of the peak current is of the order of milliampere with the velocity scale close to $c$. In this case, and at typical laboratory length scales of meters, hence, timescales of nanoseconds, the charge of the currents are of order $q\sim 10^{-12}$[C]. With these numbers, a two-orders-of-magnitude gain in the potential amplitude lifts the expected strength of the electric field in Eq. (23) from $E^{\text{sub}}_{\text{max}}\sim 10^{-3}\times 10^{2}\,[\text{V/m}]\sim 10^{-1}$ [V/m] to $E^{\text{sup}}_{\text{max}}\sim 10$ [V/m], while the velocity kick on a single electron increases from $\Delta v_{\text{sub}}\sim 10$ [m/s] to $\Delta v_{\text{sup}}\sim 10^{3}$ [m/s]. Note that at a given length scale, both the velocity kick, as well as the maximal electric field value, primarily depend on the scale of the emitted current, whose maximization should therefore be prioritized.
@@
In general, a decrease of the phase velocity of light can either be realized within particular media, as explicitly presented in section II in the case of a simple dielectric medium, or in the vacuum within adequate slow-wave structures, such as traveling-wave tubes (TWT). Of course, the conception of a realistic experimental setup to measure electromagnetic memory through the Lorentz breaking enhancement identified in this work still requires at least two significant steps that we want to address in future work. The first is a refined, quantitative computation of the memory amplitude within the given situation, taking into account finite-size- as well as expected shock-wave-effects that cannot be captured through the Green’s function of the wave equation  [^6] The computational techniques developed in the context of Cherenkov radiation might be of great use.. The second is an elaboration on the measurement technique, where it must in particular be ensured that the measurement time-scale is quick enough to capture the memory signal within the unbounded state of the emitted charges and that the memory kick can be discriminated from possible Poynting vector kicks associated with radiation pressure.
@@ IV Conclusion
While the latter case necessarily requires a reduction of the effective phase velocity of light $v_{\text{em}}<c$, the former can also be achieved at a vacuum speed of light $v_{\text{em}}=c$ with the emission of ultrarelativistic charges of $v_{q}\gtrsim 0.99c$.
@@
Although the actual divergence in the memory amplitude as well as the emitted radiative energy remains intangible with the treatment of this work, the enhancement in the vicinity of the critical directions is physical, as shown by the presence of the entirely well-defined effect in the subluminal vacuum case, and indicates the formation of a shock-wave. A crucial future task will be to compute the realistic peak in memory offset increase in specific local Lorentz breaking setups and conceptualize concrete measurement strategies of the electromagnetic memory amplitude by addressing the shock nature of the phenomenon, as well as associated finite-size and possible dispersion and absorption effects  [^7] Note, however, that the consideration of a finite acceleration time of the emitted charges has no influence on the soft limit of the radiation and only serves as a high-frequency cutoff as shown in Appendix B..
On the theory side, the work opens a novel research direction within the context of the infrared triangle [25, 26, 27, 28, 21]. It raises the question of whether the enhancement of electromagnetic memory can be formulated in terms of Lorentz-violating extensions of the leading Weinberg photon soft theorem [20], with corresponding implications on the asymptotic $U(1)$ symmetry. Moreover, since the enhancement arises from the universal Green’s function factor of the wave equation, the same mechanism is equally expected to be present in hereditary tail effects linked to sub-leading soft photon factors [44, 45, 46] and their well-known gravitational analogs [5, 47, 48] [^8] Concretely, as the tail effects are tightly connected to the solution of the wave equation as well [55, 46], the amplifying Green’s function factor in Eq. (12) will also appear in the tail contributions, as can for instance explicitly be seen in Eq.(12) of Ref. [48]..